Manuel Kauers Hagenberg, 11. April 2003 RISC-Linz

Subclauses IV

Relative Clauses and Indirect Question Clauses

Recall: Indirect Question Clauses

• Indirect question clauses are subclauses.

• They start with a question word.

• The most frequent question words are

wer, was, wo, wann, wie, warum, welches, weshalb, …

• More question words can be build by wo (+ r) + preposition.

• The question words wer and welche are declinated. The forms of wer are as follows, they don’t differ in genter and number:

case  form
Nominativ wer
Genitiv  wessen
Dativ  wem
Akkusativ wen

The forms of welches are sensitive with respect to both genter, and number:

singular  plural
case male female neutral male female neutral
Nominativ welcher welche welches welche welche welche
Genitiv  welchen welcher welches welcher welcher welcher
Dativ  welchem welcher welchem welchen welchen welchen
Akusativ welchen welche welches welche welche welche

• Examples:

Er fragt mich, welcher Bus nach Linz fährt. Er fragt mich, welchen Bus er nehmen muß. Er fragt mich, mit welchem Bus er nach Linz fahren kann. Er fragt mich, welche Busse nach Linz fahren. Er fragt mich, mit welchen Bussen er nach Linz fahren kann.

Exercise

• The following holes are supposed to be filled in by an appropriate form of

wer.

der Stift – pencil einladen – invite mitnehmen – to take with

Ich weiß nicht, Stift das ist.

Ich weiß nicht, heute kommt.

Susanne weiß, mit sie zur Uni fahren kann.

Peter weiß, er mit dem Auto zur Uni mitnehmen kann.

• Indirect question clauses are also used for reported speach, if you want to quote a question. Consider the following short dialog between you (Y) and your supervisor (S):

anmelden – to apply vorschlagen – to suggest tun – to do entscheiden – to decide

S: Wann willst Du Dich zum Rigorosum anmelden? Y: Das will ich schon im nächsten Jahr tun. Y: Welche Themen schlägst Du für das Rigorosum vor? S: Das entscheiden wir, wenn Du Dich anmeldest.

You want to report this short discussion to your colleague. Complete the following skeleton accordingly. Use indirect question clauses to convert the questions, and daß clauses to convert ordinary sentences. You can leave the wenn clause as it is.

der Betreuer – supervisor gefragt – asked geantwortet – answered

Mein Betreuer hat mich heute gefragt, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Ich habe ihm geantwortet, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Dann habe ich ihn gefragt, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ,

und er hat mir geantwortet,. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Recall: Relative Clauses

• Relative clauses serve to add information to a noun.

• The are introduced by a relative pronoun whose gender and number is according to the gender and number of the clarified noun, and whose case is according to the grammatical function in the relative clause.

• Here are the relative pronouns; they are similar to the articles.

singular  plural
Nominative der die das die die die
Gentive  dessen deren dessen deren deren deren
Dative  dem  der  dem  denen denen denen
Akkusative den  die  das  die  die  die

• Relative clauses are placed directly after the noun which is clarified.

• A relative clause clarifying a noun X may be postponed to the end of the clause in which X lives, if this simplifies the construction.

• Example:

Peter nimmt Susanne in dem Auto, das ihm gehört, mit. Peter nimmt Susanne in dem Auto mit, das ihm gehört.

Exercise

• Fill in the relative pronouns in the following sentences.

die Ablage – phys. maibox das Sekretariat – secretarie’s room gehören – belong to Symbolisches Rechnen – symbolic computation forschen – doing research die Hauptstelle – “main location” mehrere – several das Dorf – village sondern – but eigenes – own

Ein Auto, Schlüssel in der Ablage vor dem Sekretariat liegt, gehört dem Institut. Das Auto gehört dem Institut,

über Symbolisches Rechnen forscht. Das Institut gehört

zur Universität, Hauptstelle in Linz ist. Mehrere Professoren, im Institut arbeiten, wohnen in Linz. Sie kommen aber nicht mit dem Auto, dem Institut gehört, in das Dorf, in das Institut ist, sondern sie haben ihr eigenes Auto. Das Auto des Instituts ist für die Studenten da,

kein eigenes Auto haben.

What’s the difference?

• There is no formal distiction between indirect question clauses and relative clauses in English.

• What is their difference in German?

• Indirect question clauses are building blocks of the surrounding clause. There is nothing in the surrounding clause to which they refer.

• Relative clauses are no building blocks themselves. They are part of the building block of the noun or pronoun to which they refer.

• Example:

eigen – own

benutzen – to use

Wer kein eigenes Auto hat, kann das RISC-Auto benutzen. Studenten, die kein eigenes Auto haben, können das RISC-Auto benutzen.

• Translate the following sentences to German. Pay special attention to the distinction between question clauses and relative clauses.

know – wissen

anybody – jemand to look for – suchen key – der Schlüssel belong to – gehören (+ Dat) show – zeigen over there – dort anyway – sowieso could go – könnte gehen bank – die Bank

Do you know whose key this is? Do you know anybody who is looking for his key? — I don’t know to whom the key belongs which you are showing me. Bring it to Betina whose office is over there. She is anyway asking who could go to the bank.